# Classifying locally distinguishable sets: No activation across bipartitions

### v1

@@ Abstract
A set of orthogonal quantum states is said to be locally indistinguishable if they cannot be perfectly distinguished by local operations and classical communication (LOCC). Otherwise, the states are locally distinguishable. However, locally indistinguishable states may find applications in information processing protocols. In this sense, locally indistinguishable states are useful. On the other hand, it is usual to consider that locally distinguishable states are useless. Nevertheless, recent works suggest that locally distinguishable states should be given due consideration as in certain situations these states can be converted to locally indistinguishable states under orthogonality-preserving LOCC (OP-LOCC). Such a counterintuitive phenomenon motivates us to ask when the aforesaid conversion is possible and when it is not. In this work, we provide different structures of locally distinguishable product and entangled states which do not allow the aforesaid conversion. We also provide certain structures of locally distinguishable states which allow the aforesaid conversion. In this way, we classify the locally distinguishable sets by introducing hierarchies among them. In a multipartite system, this study becomes more involved as there exist multipartite locally distinguishable sets which cannot be converted to locally indistinguishable sets by OP-LOCC across any bipartition. We say this as “no activation across bi-partitions".
Multipartite entanglement, strongly local class, local indistinguishability, activation of nonlocality
@@ I Introduction
Non-local properties of quantum systems have a class exclusive from Bell nonlocality Bell nonlocality . Specifically, when a set of orthogonal quantum states cannot be perfectly distinguished by local operations and classical communication (LOCC), it reflects a fundamental nonlocal feature of quantum physics BennettPB1999 . Local distinguishability of quantum states refers to the task of identifying a state from a set of prespecified orthogonal states shared among parties separated by arbitrary distances and LOCC being the only legit class of operation BennettUPB1999 ; Walgate2000 ; Virmani ; Ghosh2001 ; Groisman ; Walgate2002 ; Divincinzo ; Horodecki2003 ; Fan2004 ; Ghosh2004 ; Nathanson2005 ; Watrous2005 ; Niset2006 ; Ye2007 ; Fan2007 ; Runyo2007 ; somsubhro2009 ; Feng2009 ; Runyo2010 ; Yu2012 ; Yang2013 ; Zhang2014 ; somsubhro2010 ; yu2014 ; somsubhro2014 ; somsubhro2016 ; bennett1996 ; popescu2001 ; xin2008 ; somsubhro2009(1) . The non-locality of orthogonal quantum states can be used for various practical purposes such as data hiding terhal58 ; divincenzo580 ; lamidatahiding ; terhaldatahiding ; chaves2020 ; wehner2020 ; winterdatahiding ; haydendatahiding , quantum secret sharing rahaman330 ; markham309 ; wang320 , and similar applications. Consequently, in the past two decades, considerable attention has been paid to the study of local distinguishability of orthogonal quantum states and the exploration of the relationship between quantum non-locality and entanglement Zhang2015 ; Wang2015 ; Chen2015 ; Yang2015 ; Zhang2016 ; Xu2016(2) ; Zhang2016(1) ; Xu2016(1) ; Halder2019strong nonlocality ; Halder2019peres set ; Xzhang2017 ; Xu2017 ; Wang2017 ; Cohen2008 ; somsubhro2018 ; zhang2018 ; Halder2018 ; Yuan2020 ; Rout2019 ; bhunia2020 ; bhunia2023 ; biswas2023 ; Zhang2019 ; bhunia2022 .
@@
Local distinguishability of quantum states plays an important role in studying the restrictions of LOCC. In 2000 Walgate et al. Walgate2000 have evinced that any two orthogonal multipartite pure states can be perfectly distinguished by allowing LOCC. Nevertheless, if there are more than two orthogonal pure states, then there can be local indistinguishability. The local indistinguishability of a set of pairwise orthogonal multipartite states is a signature of non-locality shown by those states. Since entanglement is intrinsically connected to non-locality, one can assume that mutually orthogonal product states can be perfectly distinguished by LOCC. However, entanglement is not necessary for local indistinguishability of quantum states BennettPB1999 ; BennettUPB1999 ; Zhang2015 ; Wang2015 ; Chen2015 ; Yang2015 ; Zhang2016 ; Xu2016(2) ; Zhang2016(1) ; Xu2016(1) ; bhunia2023 ; biswas2023 ; Halder2019strong nonlocality ; Halder2019peres set ; Xzhang2017 ; Xu2017 ; Wang2017 ; Cohen2008 ; Zhang2019 ; somsubhro2018 ; zhang2018 ; Halder2018 ; Yuan2020 ; Rout2019 ; bhunia2020 ; bhunia2022 ; bhuniaubb2024 ; indra2025 ; subrata2024 ; bhunia2025 . In 1999 Bennett et al. BennettPB1999 first exhibited a set of nine pure product states in a two-qutrit system, which cannot be perfectly distinguished by LOCC and presented the phenomenon of “non-locality without entanglement”. The result indicates that entanglement is not a requisite factor of local indistinguishability of quantum states. This manifests that the absence of entanglement is not sufficient to ensure the local accessibility of information. Furthermore, there is an incomplete basis for demonstrating the phenomenon of non-locality without entanglement, commonly known as the unextendible product basis (UPB). It is defined by a set of mutually orthogonal product states satisfying the condition that the orthogonal complement of the subspace, spanned by all these product states, contains no product states, i.e., this set of states cannot be extended to a complete basis by adding product states to it while preserving the orthogonality of the set Divincinzo . UPB cannot be accurately distinguished by LOCC LOCC , and the normalised projector onto the orthogonal complement of it is a mixed state which uncovers a captivating phenomenon known as bound entanglement BennettUPB1999 ; Divincinzo . Thus, these states are of considerable interest in quantum information theory.
@@
Due to practical applications, local indistinguishability of quantum states can be considered as a resource in quantum information processing. If there are only locally distinguishable sets at hands, how can we transfer them into resources that have applications in data hiding? This is what the authors of the paper bandhyopadhyay201 recently studied. In fact, they studied the following problem: is there any set of orthogonal states which can be locally distinguishable, but under an orthogonality-preserving local measurement, each outcome will lead to a locally indistinguishable set. As there are some trivial sets with this property, they introduced the concept of local irredundancy. An orthogonal set is said to be locally redundant if it remains orthogonal after discarding one or more subsystems. Otherwise, it is said to be locally irredundant. If a locally irredundant set satisfies the aforementioned property, then we say that its nonlocality can be activated genuinely, i.e., hidden nonlocality can be revealed. In Ref. bandhyopadhyay201 , the authors provided several examples of such sets with entanglement. However, deeper research on this property remains to be explored. For example, the following questions are required to be studied. Is there any multipartite locally distinguishable sets (with or without entanglement) whose nonlocality cannot be activated even if (specific) joint operations are allowed? In which multipartite state spaces can locally distinguishable sets be constructed? Answering such questions are particularly important to understand when one can have activation of nonlocality. See also subrata2024 in this regard.
In the process of studying the aforesaid questions, here we manage to construct sets of multipartite states which is not activable in any bipartition. In other words, such locally distinguishable sets cannot be transformed to a locally indistinguishable set in any bipartition under orthogonality-preserving LOCC.
@@
This is, in fact, the worst-case scenario in view of nonlocality activation. Because if we consider all subsystems together in a single location, then, anyway, there will be no local indistinguishability as we are dealing with orthogonal states here. The discovery of this class of sets also leads to a hierarchy among the multipartite locally distinguishable sets. The structures that we provide here can be easily generalised. In particular, for bipartite systems, we consider higher-dimensional Hilbert spaces compared to some known results of two-qubit or qubit-qudit cases bandhyopadhyay201 . Then, we compare between locally distinguishable product states and entangled states. The paper is organised as follows: in Sec . II, necessary definitions and other preliminary concepts are presented. In Sec. III, we provide activable and non-activable sets of product states in bipartite as well as in multipartite scenarios. In Sec . IV, we consider entangled states and present comparisons between product states and entangled states. Finally, the conclusion is drawn in Sec . V.
@@ II Preliminaries
Walgate2002 ; Halder2018 If all the POVM elements of a measurement structure, corresponding to a discrimination task of a given set of states, are proportional to the identity matrix, then such a measurement is not useful to extract information for this task and is called a trivial measurement. Conversely, should at least one POVM element not be proportional to the identity matrix, the measurement is then classified as non-trivial.
@@
Walgate2002 ; Halder2018 Consider a local measurement to distinguish a fixed set of pairwise orthogonal quantum states. Should the post-measurement states likewise exhibit the property of pairwise orthogonality, then such a measurement shall be termed an orthogonality-preserving local measurement (OPLM).
@@
Halder2019strong nonlocality A set of orthogonal quantum states is locally irreducible if it is not possible to eliminate one or more quantum states from the set by nontrivial orthogonality-preserving local measurements.
@@
A set of orthogonal quantum states is said to be locally indistinguishable if, whilst it may be possible to eliminate one or more states from the set via an OPLM, it proves impossible to completely distinguish the entire set using a non-trivial OPLM.
@@
A locally distinguishable set of multipartite orthogonal states, S, is deemed to possess hidden nonlocality of type-1 if, upon spatial separation of all constituent parties, the set may be activated by means of LOCC. We denote this by $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S})\neq 0$. Also, if a locally distinguishable set of multipartite orthogonal states $\mathcal{S}$ is said to have hidden nonlocality of type-$k$, if to activate the set by LOCC, at least $k$ parties are needed to come together, whereas all other parties are spatially separated. We denote this by $\mathcal{H}^{\text{LOCC}}_{k}(\mathcal{S})\neq 0$.
The maximum value of $k$ in $\mathcal{H}^{\text{LOCC}}_{k}(\mathcal{S})$ can be equal to $(N-1)$, because, if $k=N$, then, all parties are coming together and there is no local indistinguishability. This happens as we are dealing with orthogonal states. Naturally, in a bipartite scenario, the only case that appears is $k=1$.
@@ III Non-activable and activable product states
In this section, we first construct a class of orthogonal product states which cannot be activable by LOCC. For better understanding, we first give an example in $\mathbf{{\mathbf{C}}^{4}}\otimes\mathbf{{\mathbf{C}}^{4}}$ and then, we generalise the result.
Figure 1: Representation of product states in ${\mathbf{C}}^{4}\otimes{\mathbf{C}}^{4}$. The bottom side represents Alice’s side and top left side represents Bob’s side (this is also maintained in other figures unless explicitly stated). We represent quantum states |$\ket{i\pm\overline{i+1}}\ket{j}$ or, |$\ket{j}\ket{i\pm\overline{i+1}}$ by rectangular tiles where |$\ket{i\pm\overline{i+1}}=\frac{1}{\sqrt{2}}(\ket{i}\pm\ket{i+1}),$ for integer ‘$i$’. Each of the square tiles represent a state of the form |$\ket{j}\ket{k}$. Tile indices correspond to consecutively ordered basis states of set $\mathcal{S}_{1}$, while tile colors indicate compatible measurement setups for both parties.
Consider the set $\mathcal{S}_{1}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}\in\mathbf{\mathbf{C}^{4}\otimes\mathbf{C}^{4}}$, by,
@@
where, $|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{A},\;|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{B},$ and $|{\mathcal{\xi}}_{k}\rangle_{A}=|k\rangle_{A},\;|{\mathcal{X}}_{k}\rangle_{B}=|k\rangle_{B}$, see Fig. 1.
@@
The set $\mathcal{S}_{1}$ does not possess any activable nonlocality under orthogonality-preserving LOCC, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{1})=0.$
@@
Suppose Alice goes first, and let ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}=[m^{a}_{ij}]_{4\times 4}$ denote any arbitrary POVM operator of Alice with outcome $m$ such that the post-measurement     states $\quad\left\{\mathcal{M}_{A}^{m}\otimes I_{B}\left|\psi_{i}\right\rangle,\right.$ $|\psi_{i}\rangle\in\mathcal{S}_{1}\}$ should be mutually orthogonal. Because $m^{a}_{ij}=0$ is necessary and sufficient for $m^{a}_{ji}=0,i<j$, we will only show $m^{a}_{ij}=0,i<j$, in the following. Then, considering the states $|0\rangle_{A}|{\mathcal{X}}_{01}^{+}\rangle_{B}$ and $|1\rangle_{A}|{\mathcal{X}}_{12}^{+}\rangle_{B},$ we know $\left\langle 0\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1\right\rangle_{A}\left\langle 0+1|1+2\right\rangle_{B}=0.$ Thus, $m^{a}_{01}=m^{a}_{10}=0.$ In the same way, for the states $\{|0\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B},|2\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B}\},$ and $\{|0\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B},|{\mathcal{\xi}}_{3}\rangle_{A}|2\rangle_{B}\}$, we can compute $m^{a}_{02}=m^{a}_{20}=0,m^{a}_{03}=m^{a}_{30}=0$, respectively. Similarly if we choose the states $\{|1\rangle_{A}|{\mathcal{X}}_{12}^{+}\rangle_{B},|2\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B}\},$ and $\{|1\rangle_{A}|{\mathcal{X}}_{12}^{+}\rangle_{B},|{\mathcal{\xi}}_{3}\rangle_{A}|2\rangle_{B}\}$, we can see $m^{a}_{12}=m^{a}_{21}=0,m^{a}_{13}=m^{a}_{31}=0$, respectively. Now considering the states $\{|{\mathcal{\xi}}_{3}\rangle_{A}|2\rangle_{B},|2\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B}\}$, we have $\left\langle 2\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|3\right\rangle_{A}\left\langle 2|2+3\right\rangle_{B}=0.$ Which imply $m^{a}_{23}=m^{a}_{32}=0.$ Therefore, ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}$ is diagonal and ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}=\operatorname{diag}\left(\delta_{0},\delta_{1},\delta_{2},\delta_{3}\right)$.

### v2

# arXiv:2507.10698v2[quant-ph] 01 Mar 2026

@@ Abstract
A set of orthogonal quantum states is said to be locally indistinguishable if they cannot be perfectly distinguished by local operations and classical communication (LOCC). Otherwise, the states are locally distinguishable. Interestingly, locally indistinguishable states can have productive applications in quantum information processing protocols. In this sense, locally indistinguishable states are useful. On the other hand, it is usual to consider that locally distinguishable states are useless. Nevertheless, recent works suggest that locally distinguishable states should be given due consideration as in certain situations these states can be converted to locally indistinguishable states under orthogonality-preserving LOCC (OP-LOCC). Such a counterintuitive phenomenon motivates us to ask when the aforesaid conversion is possible and when it is not. In this work, we provide different structures of locally distinguishable product and entangled states which do not allow the aforesaid conversion. We also provide certain structures of locally distinguishable states which allow the aforesaid conversion. In this way, we classify the locally distinguishable sets by introducing hierarchies among them. In a multipartite system, this study becomes more involved as there exist multipartite locally distinguishable sets which cannot be converted to locally indistinguishable sets by OP-LOCC across any bipartition. We say this as “no activation across bi-partitions".
Multipartite entanglement, strongly local class, local indistinguishability, activation of non-locality
@@ I Introduction
Non-local properties of quantum systems have a class exclusive from Bell non-locality Bell non-locality . Specifically, when a set of orthogonal quantum states cannot be perfectly distinguished by local operations and classical communication (LOCC), it reflects a fundamental nonlocal feature of quantum physics BennettPB1999 . Local distinguishability of quantum states refers to the task of identifying a state from a set of prespecified orthogonal states shared among parties separated by arbitrary distances and LOCC being the only legit class of operation BennettUPB1999 ; Walgate2000 ; Virmani ; Ghosh2001 ; Groisman ; Walgate2002 ; Divincinzo ; Horodecki2003 ; Fan2004 ; Ghosh2004 ; Nathanson2005 ; Watrous2005 ; Niset2006 ; Ye2007 ; Fan2007 ; Runyo2007 ; somsubhro2009 ; Feng2009 ; Runyo2010 ; Yu2012 ; Yang2013 ; Zhang2014 ; somsubhro2010 ; yu2014 ; somsubhro2014 ; somsubhro2016 ; bennett1996 ; popescu2001 ; xin2008 ; somsubhro2009(1) . The non-locality of orthogonal quantum states can be used for various practical purposes such as data hiding terhal58 ; divincenzo580 ; lamidatahiding ; terhaldatahiding ; chaves2020 ; wehner2020 ; winterdatahiding ; haydendatahiding , quantum secret sharing rahaman330 ; markham309 ; wang320 , and similar applications. Consequently, in the past two decades, considerable attention has been paid to the study of local distinguishability of orthogonal quantum states and the exploration of the relationship between quantum non-locality and entanglement Zhang2015 ; Wang2015 ; Chen2015 ; Yang2015 ; Zhang2016 ; Xu2016(2) ; Zhang2016(1) ; Xu2016(1) ; Halder2019strong non-locality ; Halder2019peres set ; Xzhang2017 ; Xu2017 ; Wang2017 ; Cohen2008 ; somsubhro2018 ; zhang2018 ; Halder2018 ; Yuan2020 ; Rout2019 ; bhunia2020 ; bhunia2023 ; biswas2023 ; Zhang2019 ; bhunia2022 .
@@
Local distinguishability of quantum states plays an important role in studying the restrictions of LOCC. In 2000 Walgate et al. Walgate2000 have evinced that any two orthogonal multipartite pure states can be perfectly distinguished by allowing LOCC. Nevertheless, if there are more than two orthogonal pure states, then there can be local indistinguishability. The local indistinguishability of a set of pairwise orthogonal multipartite states is a signature of non-locality shown by those states. Since entanglement is intrinsically connected to non-locality, one can assume that mutually orthogonal product states can be perfectly distinguished by LOCC. However, entanglement is not necessary for local indistinguishability of quantum states BennettPB1999 ; BennettUPB1999 ; Zhang2015 ; Wang2015 ; Chen2015 ; Yang2015 ; Zhang2016 ; Xu2016(2) ; Zhang2016(1) ; Xu2016(1) ; bhunia2023 ; biswas2023 ; Halder2019strong non-locality ; Halder2019peres set ; Xzhang2017 ; Xu2017 ; Wang2017 ; Cohen2008 ; Zhang2019 ; somsubhro2018 ; zhang2018 ; Halder2018 ; Yuan2020 ; Rout2019 ; bhunia2020 ; bhunia2022 ; bhuniaubb2024 ; indra2025 ; subrata2024 ; bhunia2025 . In 1999 Bennett et al. BennettPB1999 first exhibited a set of nine product states in a two-qutrit system, which cannot be perfectly distinguished by LOCC and presented the phenomenon of “non-locality without entanglement”.
@@
Due to practical applications, local indistinguishability of quantum states can be considered as a resource in quantum information processing. Now, if there are only locally distinguishable sets at hands, how can we transfer them into resources that have applications in data hiding? This is what the authors of the paper bandhyopadhyay201 recently studied. In fact, they studied the following problem: is there any set of orthogonal states which can be locally distinguishable, but under an orthogonality-preserving (OP) local measurement, each outcome will lead to a locally indistinguishable set. As there are some trivial sets with this property, they introduced the concept of local irredundancy. An orthogonal set is said to be locally redundant if it remains orthogonal after discarding one or more subsystems. Otherwise, it is said to be locally irredundant. If a locally irredundant set satisfies the aforementioned property, then we say that its non-locality can be activated genuinely, i.e., hidden non-locality can be revealed. In Ref. bandhyopadhyay201 , the authors provided several examples of such sets with entanglement. However, deeper research on this property remains to be explored. For example, the following questions are required to be studied. Is there any multipartite locally distinguishable set (with or without entanglement) whose non-locality cannot be activated even if (specific) joint operations are allowed? In which multipartite state spaces can such locally distinguishable sets be constructed? Answering such questions are particularly important to understand when one can have activation of non-locality. See also subrata2024 in this regard. In the following, we provide a table (Table 1) where different activation results and corresponding references are summarized. From this table, it can be understood what has been done so far and what are the present contributions.
Results Status
Locally distinguishable sets of orthogonal entangled states can Bipartite and multipartite examples
be converted to locally indistinguishable sets under OP-LOCC. are constructed in bandhyopadhyay201 .
Similar results as described in the above row are observed Bipartite and multipartite examples
considering orthogonal product states. are constructed in Li2022 .
Locally distinguishable sets of orthogonal product states can Bipartite examples are not possible.
be converted to locally indistinguishable sets under OP-LOCC, Thus, multipartite examples are
such that the sets exhibit local indistinguishability across constructed in GhoshStrongActivation2022 .
every bipartition.
Locally distinguishable sets of orthogonal states cannot be Bipartite examples are not possible.
converted to locally indistinguishable sets under OP-LOCC Thus, multipartite examples are
when all parties are parties are spatially separated. However, constructed in subrata2024 .
it can be done when two parties collaborate and perform joint
measurements.
There are instances when locally distinguishable sets of Bipartite examples are not possible.
orthogonal states cannot be converted to locally indistinguishable Thus, multipartite examples are
sets under OP-LOCC even if joint measurements are allowed (all constructed here. We also develop
but one parties are allowed to collaborate and perform joint methodology for different constructions.
measurements). A possible classification among existing
and presently provided structures are
discussed. (present contributions)
Table 1: Summary of LOCC activation results and corresponding references are given.
@@
In the process of studying the aforesaid questions, here we manage to construct sets of multipartite states which are not activable in any bipartition. In other words, such locally distinguishable sets cannot be transformed to a locally indistinguishable set in any bipartition under orthogonality-preserving LOCC. We say this as the worst-case scenario in view of non-locality activation. Because if we consider all subsystems together in a single location, then, anyway, there will be no local indistinguishability as we are dealing with orthogonal states here. The discovery of this class of sets also leads to a hierarchy among the multipartite locally distinguishable sets. The structures that we provide here can be easily generalised. In particular, for bipartite systems, we consider higher-dimensional Hilbert spaces compared to some known results of two-qubit or qubit-qudit cases bandhyopadhyay201 . Then, we compare between locally distinguishable product states and entangled states. The paper is organised as follows: in §II, necessary definitions and other preliminary concepts are presented. In §III, we provide activable and non-activable sets of product states in bipartite as well as in multipartite scenarios. In §IV, we consider entangled states and present comparisons between product states and entangled states. Finally, the conclusion is drawn in §V.
@@ II Preliminaries
Walgate2002 ; Halder2018 If all the POVM elements of a measurement structure, corresponding to a discrimination task of a given set of states, are proportional to the identity matrix, then such a measurement is not useful to extract information for this task and is called a trivial measurement. Conversely, there should be at least one POVM element not proportional to the identity matrix, then the measurement is called as non-trivial.
@@
Walgate2002 ; Halder2018 Consider a local measurement to distinguish a fixed set of pairwise orthogonal quantum states. If the post-measurement states likewise exhibit the property of pairwise orthogonality, then such a measurement shall be termed as an orthogonality-preserving local measurement (OPLM).
@@
Halder2019strong non-locality A set of orthogonal quantum states is locally irreducible if it is not possible to eliminate one or more quantum states from the set by nontrivial orthogonality-preserving local measurements.
@@
A set of orthogonal quantum states is said to be locally indistinguishable if, whilst it may be possible to eliminate one or more states from the set via an OPLM, it is impossible to completely distinguish the entire set using a non-trivial OPLM.
@@
A locally distinguishable set of multipartite orthogonal states, S, is deemed to possess hidden non-locality of type-1 if, upon spatial separation of all constituent parties, the set may be activated by means of LOCC. We denote this by $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S})\neq 0$. Also, if a locally distinguishable set of multipartite orthogonal states $\mathcal{S}$ is said to have hidden non-locality of type-$k$, if to activate the set by LOCC, at least $k$ parties are needed to come together, whereas all other parties are spatially separated. We denote this by $\mathcal{H}^{\text{LOCC}}_{k}(\mathcal{S})\neq 0$.
The maximum value of $k$ in $\mathcal{H}^{\text{LOCC}}_{k}(\mathcal{S})$ can be equal to (N1)(N-1), because, if $k=N$, then, all parties are coming together and there is no local indistinguishability. This happens as we are dealing with orthogonal states. Naturally, in a bipartite scenario, the only case that appears is $k=1$. In this work, when we consider a particular value of $k$, then to check activability/non-activability, we consider local distinguishability/indistinguishability in $k$-partitions only. If for k=N1k=N-1, it is not possible to convert a locally distinguishable set to an indistinguishable one under OP-LOCC, i.e., when we get non-activability in all $k$-partitions for maximum value of $k$, then we say it as worst case scenario here.
@@ III Non-activable and activable product states
In this section, we first construct a class of orthogonal product states which cannot be activable by LOCC. For better understanding, we first give an example in ${\mathbb{C}}^{4}\otimes{\mathbb{C}}^{4}$ and then, generalise the result.
[…]
Consider the set $\mathcal{S}_{1}\subseteq\mathbb{C}^{4}\otimes\mathbb{C}^{4}$, by,
@@
where, $\mathbf{|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{A},\;|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{B}},$ and $\mathbf{|{\mathcal{\xi}}_{k}\rangle_{A}=|k\rangle_{A},\;|{\mathcal{X}}_{k}\rangle_{B}=|k\rangle_{B}}$, see Fig. 1.
@@
The set $\mathcal{S}_{1}$ does not possess any activable non-locality under orthogonality-preserving LOCC, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{1})=0$.
@@
Suppose Alice makes the first measurement. Let ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}=[m^{a}_{ij}]_{4\times 4}$ denote any arbitrary POVM operator of Alice with outcome $m$ such that the post-measurement states $\left\{\mathcal{M}_{A}^{m}\otimes I_{B}\left|\psi_{i}\right\rangle,|\psi_{i}\rangle\in\mathcal{S}_{1}\right\}$ are mutually orthogonal. As, $m^{a}_{ij}=0$ is necessary and sufficient for $m^{a}_{ji}=0,\,i<j$. We only need to show that $m^{a}_{ij}=0,i<j$, in the following. For the states $\mathbf{|0\rangle_{A}|{\mathcal{X}}_{01}^{+}\rangle_{B}}$ and $\mathbf{|1\rangle_{A}|{\mathcal{X}}_{12}^{+}\rangle_{B}},$ it is easy to see that $\left\langle\mathbf{0}\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|\mathbf{1}\right\rangle_{\mathbf{A}}\mathbf{\left\langle 0+1|1+2\right\rangle_{B}}=0.$ Thus, $m^{a}_{01}=m^{a}_{10}=0.$
In the same way, for the states $\{\mathbf{|0\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B},|2\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B}}\},$ and $\{\mathbf{|0\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B},|{\mathcal{\xi}}_{3}\rangle_{A}|2\rangle_{B}}\}$, we can see that $m^{a}_{02}=m^{a}_{20}=0,m^{a}_{03}=m^{a}_{30}=0$, respectively. Similarly, if we choose the states $\{\mathbf{|1\rangle_{A}|{\mathcal{X}}_{12}^{+}\rangle_{B},|2\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B}}\},$ and $\{\mathbf{|1\rangle_{A}|{\mathcal{X}}_{12}^{+}\rangle_{B},|{\mathcal{\xi}}_{3}\rangle_{A}|2\rangle_{B}}\}$, we get $m^{a}_{12}=m^{a}_{21}=0,m^{a}_{13}=m^{a}_{31}=0$.
Now considering the states $\{\mathbf{|{\mathcal{\xi}}_{3}\rangle_{A}|2\rangle_{B},|2\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B}}\}$. For this, we have $\left\langle\mathbf{2}\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|\mathbf{3}\right\rangle_{\mathbf{A}}\mathbf{\left\langle 2|2+3\right\rangle_{B}}=0$, which implies that $m^{a}_{23}=m^{a}_{32}=0.$ Therefore, with respect to the basis $\mathcal{S}_{1}$, ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}$ is diagonal and ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}$ can be written as $\mathrm{diag}\left(\delta_{0},\delta_{1},\delta_{2},\delta_{3}\right)$.